OpenStudy (anonymous):

Solve using substitution: -y − 5x = -7 2y − 2x = 14 A) Infinitely many solutions B) No solution C) (7/5, 14) D) (0, 7)

OpenStudy (anonymous):

Solution is (0,7).. So, D is the answer...

OpenStudy (anonymous):

Find y from the first equation : y = 7 - 5x Put it in second equation: we get, x = 0 and y = 7...

OpenStudy (anonymous):

considering the given equations... equation 1 can be writen as 5x+y=7, equation 2 can be written as -2x+2y=14 we need to solve these two equations now. we can get to the ans (x,y)=(0,7)